Select one company from your portfolio (Starbucks Corporation SBUX) and determine the following for the financial period ended March 31, 2019. Note: You will need to locate the financial statements for this company from sources like: Yahoo Finance, Google Finance, or the Company’s web site. Please name your company and show the calculations:

(1). Current Ratio

(2). Debt Ratio

(3). Profit Margin

(4). Return on Assets

(5). Return on Equity
Answer #1

Company Name : Starbucks Corporation

Based on Financials as of 30/09/2018: (in thousand $)

Current Assets = 1,24,94,200, Current Liability = 56,84,200, Total Debt = 90,90,200, Equity = 11,69,500, Sales = 2,47,19,500, Net profit = 45,18,300, Total Assets = 2,41,56,400.

1) Current Ratio = (Current Assets/ Current Liabilities) = (1,24,94,200/56,84,200) = 2.19

2) Debt Ratio = (Total Debt/ (Equity+debt) = (90,90,200)/(11,69,500+90,90,200) = 88.60%

3) Profit Margin = (Net profit/ Sales) * 100 = (45,18,300/2,47,19,500) *100 = 18.27%

4) Return on Assets = (Net profit/ Total Assets) * 100 = (45,18,300)/ (2,41,56,400)*100 = 18.70%

5) Return on Equity = (Net profit/ Equity) * 100 = (45,18,300)/(11,69,500)*100 = 386.34%